OpenStudy (anonymous):

how do i find the product of this and simplify if possible? -5/9 times 6/8

OpenStudy (theedwardsfamily):

https://www.symbolab.com/

OpenStudy (kfireangel):

try making the fractions into decimals, that might help

OpenStudy (alexandervonhumboldt2):

\[-\frac{ 5 }{ 9 }*\frac{ 6 }{ 8 }=-\frac{ 5*6 }{ 9*8 }=\]